OpenStudy (anonymous):

Find the critical points (min. max.) & points of inflection of y = 1/3³ - 1/2x² - 2x Someone help me please? : )

OpenStudy (anonymous):

is that 1/3x^3 or is 1/3^3

OpenStudy (anonymous):

1/3x

OpenStudy (anonymous):

Sorry typo :D

OpenStudy (anonymous):

1/3x³

OpenStudy (anonymous):

finding x u deferentiate this gives x^2-x-2=0(this is ur y') x=2 or x=-1 sub x into the original equation y=-13/3 or y=13/6 def y'' to get inflection

OpenStudy (anonymous):

kirstinpearl: a word of advice - the way you've typed the question is confusing - 1/3x^3 can mean 'one third times x cubed' or 'one divided by 3 x cubed'. it is better to use the equation generator or draw function If you are not happy with these then type one third x cubed as (1/3)x^3 to avoid confusion. if the first option is correct, then remainder has solved it for you.

OpenStudy (anonymous):

y" = 2x-1 for x = 2, y" is positive so (2,-13/3) is a minimum for x = -1, y" is negative so (-1,13/6) is a maximum